Evaluate the ∫(x+2/x-2)dx


1
Expert's answer
2021-02-25T05:04:13-0500

"\int \frac{x+2}{x-2} dx = \int \frac{x-2+4}{x-2} dx = \int \frac{x-2}{x-2} dx + \int \frac{4}{x-2} dx = \int dx +4 \int \frac{1}{x-2} dx" "= x + 4ln\mid x-2 \mid + C"